Calculate the solubility of kcl in g/dm³ if 5gram of the salt was dissolved in 50cm³ of water at 40 degree celius


1
Expert's answer
2021-05-24T07:22:32-0400

Using the information provided and the conversion between cm3 and dm3 ("1\,dm^3= 1000\,cm^3") we calculate the solubility as:


"S = \frac{m}{V} = \frac{5\,g\,KCl}{50\,\cancel{cm^3}}*\frac{1000\,\cancel{cm^3}}{1\,dm^3}"


"S = \frac{5000\,g\,KCl}{50\,dm^3}= \frac{100\,g\,KCl}{dm^3}"


At last, we find that the solubility of KCl at 40 °C is 100 g/dm3